### v2

# No Vulnerability Data, No Problem: Towards Predicting Mean Time To Remediate In Open Source Software Dependencies

Abstract.
Timely remediation of vulnerabilities in software dependencies is critical for the security of the software supply chain. As such, researchers have proposed tools and metrics to help practitioners assess the security practices of each of their dependencies. Conceptually, a dependency-focused Mean-Time-To-Remediate (MTTR) metric can provide a historical perspective on how long it takes a given package to update vulnerable versions of its dependencies. However, existing MTTR metrics focus on a package fixing bugs in its own code, not its dependencies. Simultaneously, existing dependency update metrics do not aggregate values for the entire package and are not sensitive to aspects important for vulnerabilities (e.g., floating version constraints). The goal of this study is to aid industry practitioners, including developers, in assessing the risk of dependencies through a novel metric approximating mean-time-to-remediate vulnerabilities in their dependencies that is evaluated by an empirical study. We propose a novel algorithm for computing MTTR called MTTRdep and a companion metric called Mean-Time-To-Updatedep (MTTUdep), which considers all version updates, including vulnerability fix updates. We conduct a large-scale study using $163,207$ packages in npm, PyPI, and Cargo, of which only $22,513$ packages produce MTTRdep because of the lack of vulnerability data. We further study how package characteristics (e.g., contributors and version counts) influence MTTUdep and MTTRdep and explore how long packages retain outdated vulnerable dependencies in npm, PyPI, and Cargo. Our results indicate that industry practitioners can reliably use MTTUdep as a proxy for MTTRdep when available vulnerability data is insufficient.
1. Introduction
Vulnerable dependencies are widely present in both open-source software (OSS) and proprietary codebases. According to the Synopsys 2025 “Open Source Security and Risk Analysis Report” (blackduck, ), 86% of codebases contain at least one vulnerable open source dependency, and 81% of codebases contain high or critical risk vulnerabilities resulting from dependencies. The majority of codebases had vulnerable dependencies for more than two years despite the availability of a fixed version (ossf-scorecard, ; blackduck-2021, ). This delay occurs because of fear of breaking changes and the cost associated with updating vulnerable dependencies to fixed versions (derrKeepMeUpdated2017, ).
Industry practitioners desire metrics that help assess the trustworthiness, risk, and security practices of open-source dependencies. For example, OpenSSF Scorecard (ossf-scorecard, ) considers 18 security practices to assess a package, including the “maintained” check, which determines if the project is maintained by checking activity in the last 90 days, and the “vulnerabilities” check detects if there are unfixed known vulnerabilities in the project or its dependencies.
Mean-Time-To-Remediate (MTTR) is also often discussed in the context of vulnerability management (sonatype-2019, ). In contrast to Scorecard’s ‘maintained’ metrics, MTTR provides a historical perspective. However, MTTR has largely been discussed in the context of a project fixing its own vulnerabilities, not vulnerabilities in dependencies (gokhaleTimeStructureBased1999, ; calvoApplyingGoalQuestion2023, ; johnsonSurveySoftwareTools1988, ; morrisonMappingFieldSoftware2018, ). Dependency update metrics such as technical lag (gonzalez-barahona_technical_2017, ) consider management of dependency versions; however, they have not been applied to vulnerabilities. Conceptually, a lower value in a dependency update metric suggests faster and more consistent updates, signaling ongoing maintenance and reliability (sonatype-2024, ), which could help assess the security risks of adopting a package. However, the connection between dependency updates and exposure to vulnerabilities in dependencies has not yet been demonstrated.
Practitioners assessing the risk associated with using a given dependency face two key challenges when considering how it manages vulnerabilities in its dependencies.
1. Insufficient Vulnerability Data. Currently, only 2.4K of npm packages have known vulnerabilities, while the ecosystem hosts over 3.27M packages. Practitioners can only compute MTTR for the dependents of those 2.4k packages.
2. Metric Limitations. The calculations used by existing dependency update metrics often do not handle floating version constraints or are not designed to handle vulnerabilities (coxMeasuringDependencyFreshness2015, ; gonzalez-barahona_technical_2017, ). Existing metrics also often calculate an update time for each of a dependency and for each version of a package, making it difficult to understand that package’s update practices as a whole. Finally, there are no existing publicly available dependency update metrics specifically for vulnerabilities. While some industry reports use MTTR in this way, their specific calculations are not available.
The goal of this study is to aid industry practitioners, including developers, in assessing the risk of dependencies through a novel metric approximating mean-time-to-remediate vulnerabilities in their dependencies that is evaluated by an empirical study. Specifically, we hypothesize that a package’s dependency update metric (Mean-Time-To-Update, MTTU) can serve as an indicator of how quickly the package would remediate security vulnerabilities (MTTR) in dependencies, even if the package never had a vulnerable dependency before. We also overcome challenges with dependency update calculations and propose novel algorithms for computing MTTU and MTTR, which we denote MTTUdep and MTTRdep, respectively.
We use MTTUdep and MTTRdep to answer the following research questions: RQ1:Can MTTUdep be used as a proxy for MTTRdep? RQ2:How do package characteristics influence MTTUdep and MTTRdep? RQ3:How long do packages have outdated dependencies, and how long do packages have vulnerable dependencies?
To answer these questions, we first collect package version release information and dependency relations of packages from the npm, PyPI, and Cargo ecosystems and compute MTTUdep and MTTRdep for $163,207$ packages. For RQ1, we perform an equivalence analysis using TOST (two one-side t-tests) and analyze correlations between MTTUdep and MTTRdep. For RQ1, we use correlation tests to understand the association between various package characteristics (e.g., contributors count, version count). We further conduct regression analysis to quantify which package characteristics matter more in influencing MTTUdep and MTTRdep values. Finally, for RQ1, we analyze the distributions of MTTUdep and MTTRdep to explore how long packages retain outdated and vulnerable dependencies in the ecosystems.
Contributions. In summary, this paper contributes (1) a detailed algorithm and process for quantifying the dependency update practice of a package using our novel metrics; (2) statistical hypothesis testing on using MTTUdep as a proxy for MTTRdep, when vulnerability data is not available; (3) a large-scale analysis of the dependency update metrics in npm, PyPI, and Cargo packages; (4) correlation and regression analysis illustrating which package characteristics impact the likelihood of higher MTTUdep and MTTRdep; and (5) actionable insights for open-source developers and supply chain security researchers based on our findings.
We provide our replication package in Zenodo (zenodo-artifact, ) (available for reviewers till Oct’25). Upon acceptance of the paper, we will make it public.
2. Related Work
In this section, we provide a brief overview of the related work.
Technical Lag. The concept of technical lag was first introduced by Gonzalez-Barahona et al. (gonzalez-barahonaTechnicalLagSoftware2017, ) for OSS packages. Essentially, “technical lag” quantifies how quickly software systems fall behind as new versions and updates are released. Zerouali et al. (zeroualiEmpiricalAnalysisTechnical2018, ) applied “technical lag” to the context of dependencies and conducted an empirical analysis of package dependency updates in the npm ecosystem. They found that outdated dependencies induce a median technical lag of 3.5 months in npm. Building on this, Decan et al.  (decanEvolutionTechnicalLag2018, ) conducted a longitudinal empirical study of ‘technical lag’ in the npm dependency network and explored how technical lag increases over time. They observed that technical lag for most npm packages increases during their lifespan, and technical lag occurred mainly due to the minor and patch releases of a dependency.
Further research by Zerouali et al. (zeroualiFormalFrameworkMeasuring2019, ) propose a formal framework for measuring technical lag in software ecosystems. They analyzed 4M releases of 500K npm packages considering the evolution of technical lag over time. They found that technical lag induced by direct dependencies in npm packages increases over time due to missed updates, including major releases. Stringer et al. (stringerTechnicalLagDependencies2020, ) study the technical lag of dependencies in a large-scale cross-ecosystem fashion containing packages from 14 package managers. They found that pinned dependencies are the main reason behind technical lag. Zerouali et al. (zeroualiMultidimensionalAnalysisTechnical2021, ) expand the idea of ‘technical lag’ into multiple dimensions (package lag, time lag, version lag, vulnerability lag, and bug lag) and study the technical lag in 140K Docker images. They found that the median time lag of community Docker images is over a year. Although previous studies have explored technical lag in terms of general dependency updates, understanding vulnerable dependency update practice was not their goal. In contrast, we study the dependency update metric and vulnerable dependency update metric and investigate their relationship.
Outdated and vulnerable dependencies. Kula et al. (kulaTrustingLibraryStudy2015, ) analyze the latency in adopting the latest version of a dependency in the Maven ecosystem. Their study found that developers are more likely to adopt the latest version for newly added dependencies than existing ones. In a follow-up study, Kula et al. (kula_developers_2018, ) examined library migration across GitHub projects and found that 81.5% of the projects keep using outdated dependencies. Cox et al. (coxMeasuringDependencyFreshness2015, ) introduce the concept of ‘dependency freshness’ to study dependency updatedness in the Maven ecosystem. They found that only $16.7\%$ of the dependencies display no update lag. Derr et al. (derrKeepMeUpdated2017, ) identify the root causes of outdated dependencies in the Android ecosystem and find that developers do not update their third-party library dependencies due to fear of breaking changes, lack of knowledge, and lack of motivation. Wang et al. (wangEmpiricalStudyUsages2020, ) conduct an empirical study on dependency update analysis on OSS packages and find that 50% of the packages use outdated dependencies. Huang et al. (huangCharacterizingUsagesUpdates2022, ) extend Wang et al. (wangEmpiricalStudyUsages2020, )’s study and find that one-third of the projects have a lag of one major version from the latest library version.
Pashchenko et al. (pashchenkoVulnerableOpenSource2018, ) studied the most used Java dependencies in SAP software and found that only updating the dependencies’ version can remove 81% vulnerable dependencies. Kula et al. (kulaDevelopersUpdateTheir2018, ) studied the update behavior of developers w.r.t. security advisories and found that developers do not update their vulnerable dependencies regularly. Kumar et al. (kumar_comprehensive_2024, ) conducted a study to understand how widespread vulnerabilities are and how quickly they are being fixed. They found that for most programming languages, a critical vulnerability persists on average for over a year before being fixed. Studies in outdated dependencies and vulnerable dependencies are focused on either all updates or only security updates but not both. In contrast, we focus on both outdated and vulnerable dependencies and explore their relationship using our proposed metrics.
MTTU and MTTR. The metrics MTTU and MTTR have been used in the software reliability and maintenance domain for a long time (gokhaleTimeStructureBased1999, ; calvoApplyingGoalQuestion2023, ; johnsonSurveySoftwareTools1988, ; morrisonMappingFieldSoftware2018, ). Researchers have also studied different security metrics (time to close bug/vulnerability, window of exposure, vulnerability count) (davidsonGoodBadUgly2009, ; beresUsingSecurityMetrics2009, ) of various categories (time metric, vulnerability metric) in the software security domain, but these metrics are focused on packages having the vulnerability. In our study, we focus on measuring dependency update metrics for packages having vulnerable dependencies, not the vulnerable package itself.
MTTR has also been used in different contexts in the industry, e.g., measuring the package’s security (mttr-dark-reading, ; mttr-plextrac, ) and measuring the package’s security in terms of dependency (sonatype-2019, ). The procedures for measuring MTTR and MTTU are often proprietary and not disclosed for academic research. For example, Sonatype’s 2024 report (sonatype-2024, ) measured MTTU and MTTR for Maven, but we could not reproduce it since the methodology is not available.
3. Challenges Applying Existing Update Metrics
In this section, we first provide an example of how the most prominent update metric with a published algorithm available in the literature, technical lag, works. Then, we describe the design gaps present in technical lag when measuring the updatedness of dependencies and why new metrics are needed.
Table 1. tLag measurement of codemod-cli package with one of its dependency, simple-git.
codemod-cli version: date simple-git constraint lastAllowed(simple-git) version: date latest(simple-git) version: date tLag
$0.8.6:\operatorname{2022-03-03}$ ${}^{\wedge}1.130.0$ $1.132.0:\operatorname{2020-03-12}$ $3.2.6:\operatorname{2022-02-17}$ $707$
$0.8.7:\operatorname{2022-03-05}$ ${}^{\wedge}2.48.0$ $2.48.0:\operatorname{2021-12-01}$ $3.2.6:\operatorname{2022-02-17}$ $78$
$0.9.0:\operatorname{2022-03-08}$ ${}^{\wedge}2.48.0$ $2.48.0:\operatorname{2021-12-01}$ $3.2.6:\operatorname{2022-02-17}$ $78$
$0.9.1:\operatorname{2022-03-17}$ ${}^{\wedge}2.48.0$ $2.48.0:\operatorname{2021-12-01}$ $3.3.0:\operatorname{2022-03-11}$ $100$
$0.9.2:\operatorname{2022-03-22}$ ${}^{\wedge}3.4.0$ $3.4.0:\operatorname{2022-03-18}$ $3.4.0:\operatorname{2022-03-18}$ 00
3.1. Technical Lag Using tLag
Time lag ($tLag$(zerouali_empirical_2018, ; decan_evolution_2018, ) is a way to measure technical lag to assess the outdatedness of a dependency in terms of time. Conceptually, $tLag$ measures the time difference between the release time of the latest version of a dependency and the release time of the version of the dependency used by a package, at the time the package was released.
Let $pkg$ be a package, and $pkg_{v}$ indicate specific version $v$ of $pkg$. Let $dep$ be a direct dependency of $pkg$. When specifying $dep$ as a dependency, $pkg_{v}$ specifies a dependency constraint $pkg_{v}.dep.\bar{c}$. Package managers use dependency constraints to resolve the highest version that satisfies the constraint. For example, a dependency constraint $pkg_{v}.dep.\bar{c}$ of “${}^{\wedge}1.2.3$” is satisfied for the range [¿=1.2.3, ¡ 2.0.0]. We use the function $lastAllowed(dep,\bar{c},t)$ to denote the version resolution of $dep$ for constraint $\bar{c}$ at time $t$. We further use the function $latest(dep,t)$ to denote the latest version of $dep$ at time $t$ and $time(\cdot)$ to denote the time a given version is released. Hence, the $tLag$ of $pkg_{v}$ for $dep$ is formally defined as follows:
(1) $tLag(pkg_{v},dep)=time(latest(dep,time(pkg_{v})))\\ -time(lastAllowed(dep,pkg_{v}.dep.\bar{c},time(pkg_{v})))$
Figure 1. Illustration of tLag, and MTTUdep/MTTRdep calculation using codemod-cli’s dependency relationship with simple-git. Blue time quantum indicates outdated dependency, red time quantum indicates outdated and vulnerable dependency, and yellow quantum indicates updated dependency.
Table 1 shows an example computation of technical lag for codemod-cli’s dependency on simple-git. The first two columns of Table 1 indicate the version with the date of different releases of package codemod-cli and the version constraint specified for dependency simple-git. The “lastAllowed(simple-git)” column, represents the resolved version of dependency simple-git with specified constraint at the time of the release of codemod-cli. For computing the technical lag, we need to know the latest available version of the dependency simple-git at the time of the release of each version of codemod-cli according to Eq. 1, which represents in column “latest(simple-git)”. The tLag column is the computed technical lag. We subtract the release date of lastAllowed(simple-git) from the release date of latest(simple-git) to obtain the technical lag for each release of codemod-cli. The computed tLag of five versions of codemod-cli were 707, 78, 78, 100, and 0, consecutively, in Table 1 and Fig. 1.
3.2. Design Gaps
In this section, we analyze the gaps available in tLag, which are used as motivators of the design choice of our proposed metrics.
1
Handling Floating Version Constraints. Floating version constraints are available in all major OSS ecosystems and are considered a good practice (he_pinning_2025, ; jafari_dependency_2023, ), as a package gets automatic updates whenever a newer version of the dependency is released. However, tLag does not sufficiently account for automatic updates that are allowed by floating version constraints. For example, tLag only checks the latest available version of the dependency against the package’s used dependency version when the package releases a new version. Even if the package releases a new version, and after that, the dependency releases a newer version that can be auto-updated by the package constraint, tLag cannot model this case in its design. Accounting for floating labels at the time of a vulnerability fix is particularly important when considering exposure to vulnerabilities, as an auto-update may automatically remediate a vulnerability.
2
Package-Level Metric. A dependency update metric should provide a way to aggregate all data points into one data for each package so that developers can make sense of the metrics and use them for comparison if desired. One of the findings of Wermke et al. (wermkeAlwaysContributeBack2023, ) was that developers’ dependency selection metrics and criteria focus on quickly accessible numbers and facts, such as downloads, GitHub stars, and time since last release, substantiating the necessity of having easily comprehensible metrics. However, tLag computes the technical lag for each version of the package for a dependency as shown in Table 1. For developers, understanding what to do with all of these tLag values for a package with multiple releases and multiple dependencies is hard. Having an aggregation would improve the understandability of dependency update metrics.
3
Handling Vulnerabilities. A security-oriented dependency update metric should incorporate the update characteristics of vulnerable dependencies. Technically, tLag’s measurement can be modified to only consider the dependency’s security fixes adopted by the package in its formal framework. However, no previous work measured tLag for measuring vulnerable dependency update practices.
4. Proposed Dependency Update Metrics
In this section, we first describe our design choices that fill the gaps presented in Section 3.2 for dependency update metrics measurement. Then, we provide an example of the nominal case using our metrics. We then provide a detailed methodology for defining the dependency update metrics in Section 4.3.
4.1. High-Level Overview
We design our metrics to only consider direct dependencies since the package only has control over which version to use for the direct dependencies. We begin by examining each of the $<$package, dependency$>$ relationships from a temporal perspective. We then split each $<$package, dependency$>$ relations into multiple intervals based on when the package or the dependency has a newer release (major, minor, or patch release). Our decision to split into intervals is to capture the benefits of floating version constraints without incentivizing this practice (Gap 1). [^1] To the best of our knowledge, no academic work quantifies optimal dependency specification (e.g., pinning vs floating version constraints) for balancing security benefits and the cost of maintenance. Therefore, our design does not discriminate between ways to specify dependency constraints (he_pinning_2025, ). Then, at each interval, the dependency constraint set by the package for the dependency is resolved with only the versions available at the beginning of the interval. We do this to make sure that the dependency resolution accounts for the historical version releases of the dependency.
After splitting into intervals, we mark “updated” as ‘false’ if the resolved dependency version for that interval does not match the highest available version of the dependency at that time and ‘true’ otherwise. Similarly, we mark “remediated” as ‘false’ if the resolved dependency version is vulnerable with a fixed version being available at the beginning of that interval and ‘true’ otherwise (Gap 3). We then aggregate the “updated” and “remediated” information of each $<$package, dependency$>$ to compute the MTTUdep and MTTRdep of each package (Gap 2).
4.2. Nominal Case With Our Metrics
Table 2. Running example of codemod-cli package with one of it’s depndency simple-git.
row pkg pkg version dep dep constraint dep version dep highest rel. Interval start Interval end updated remediated
1 codemod-cli $0.4.0$ simple-git ${}^{\wedge}1.130.0$ $1.132.0$ $2.13.1$ $\operatorname{2020-07-16}$ $\operatorname{2020-07-17}$ false true
$\ldots$ $\ldots$ $\ldots$ $\ldots$ $\ldots$ $\ldots$ $\ldots$ $\ldots$ $\ldots$ $\ldots$
82 codemod-cli $0.8.6$ simple-git ${}^{\wedge}1.130.0$ $1.132.0$ $3.2.6$ $\operatorname{2022-03-03}$ $\operatorname{2022-03-05}$ false true
83 codemod-cli $0.8.7$ simple-git ${}^{\wedge}2.48.0$ $2.48.0$ $3.2.6$ $\operatorname{2022-03-05}$ $\operatorname{2022-03-08}$ false true
84 codemod-cli $0.9.0$ simple-git ${}^{\wedge}2.48.0$ $2.48.0$ $3.2.6$ $\operatorname{2022-03-08}$ $\operatorname{2022-03-11}$ false true
85 codemod-cli $0.9.0$ simple-git ${}^{\wedge}2.48.0$ $2.48.0$ $3.3.0$ $\operatorname{2022-03-11}$ $\operatorname{2022-03-17}$ false false
86 codemod-cli $0.9.1$ simple-git ${}^{\wedge}2.48.0$ $2.48.0$ $3.3.0$ $\operatorname{2022-03-17}$ $\operatorname{2022-03-18}$ false false
87 codemod-cli $0.9.1$ simple-git ${}^{\wedge}2.48.0$ $2.48.0$ $3.4.0$ $\operatorname{2022-03-18}$ $\operatorname{2022-03-22}$ false false
88 codemod-cli $0.9.2$ simple-git ${}^{\wedge}3.4.0$ $3.4.0$ $3.4.0$ $\operatorname{2022-03-22}$ $\operatorname{2022-03-29}$ true true
$\ldots$ $\ldots$ $\ldots$ $\ldots$ $\ldots$ $\ldots$ $\ldots$ $\ldots$ $\ldots$ $\ldots$ $\ldots$
In this section, we explain how to calculate the dependency update metrics. As a running example, we consider the dependency relation between codemod-cli and simple-git, as shown in Table 2.
We split the dependency relations into multiple intervals based on the release of a new version of the package “codemod-cli” or dependency “simple-git”. In each interval, for resolving the dependency constraint by the package codemod-cli, we only consider the dependency versions available at the beginning of the interval (“interval start”). From $V0.4.0$ to $V0.8.6$ (row 1-82), codemod-cli has the dependency constraint simple-git ${}^{\wedge}1.130.0$ which resolves into $V1.132.0$. simple-git has the highest release $V2.13.1$ at $\operatorname{2020-07-16}$ (row 1) and $V3.2.6$ at $\operatorname{2022-03-03}$. As a result, codemod-cli does not have the highest available release of simple-git in these intervals (rows 1-82). In $V0.8.7$, codemod-cli updated the constraint for simple-git to ${}^{\wedge}2.48.0$. Even with this update, codemod-cli has not changed the constraint of simple-git to the available highest major version 3 and rather stayed at major version 2. codemod-cli have ${}^{\wedge}2.48.0$ constraint for simple-git from $V0.8.7$ to $V0.9.1$. The false value in the “updated” column means that the package has an outdated version of the dependency in this interval.
At the beginning of row 85, the resolved version of simple-git $V2.48.0$ was found vulnerable to four vulnerabilities (CVE-2022-24066 (CVE-2022-24066, ), CVE-2022-24433 (CVE-2022-24433, ), CVE-2022-25912 (CVE-2022-25912, ), and CVE-2022-25860 (CVE-2022-25860, )). In this example, we only consider CVE-2022-24433 as a vulnerability, which was fixed in $V3.3.0$, released on $\operatorname{2022-03-11}$. For this reason, we mark the intervals from rows 85-87 as “remediated” = false since codemod-cli has a vulnerable version of simple-git even though a fixed version is available. The false value in the “remediated” column means that the package has an outdated and vulnerable version of the dependency in this interval. In $V0.9.2$, codemod-cli bumped the constraint of simple-git to ${}^{\wedge}3.4.0$ which resulted in $V3.4.0$, a fixed version of the above vulnerability. So codemod-cli has the highest available version of simple-git in this interval (row 88) and so marked as true in the “remediated” column.
When computing the Time-To-Update (TTUdep) for dependency relation between codemod-cli and simple-git, we sum up the intervals marked false in the “updated” column. For codemod-cli, the TTUdep is the delta between $\operatorname{2020-07-16}$ to $\operatorname{2022-03-22}$ (rows 1 - 87) which is $614$ days. Similarly, for computing Time-To-Remediate (TTRdep), we sum up the intervals marked false in the “remediated” column. For codemod-cli, the TTRdep is the delta between $\operatorname{2022-03-11}$ to $\operatorname{2022-03-22}$ (rows 85 - 87) which is $11$ days.
4.3. Metrics Definitions
We present a formal definition of our proposed metrics in this section.
Metric 1 ($\operatorname{\mathbf{Mean-Time-To-Update}}_{dep}$ : $\mathbf{MTTU}_{dep}$).
MTTUdep of a package is the average aggregated time the package uses an outdated direct dependency version in its lifetime.
We have described how to calculate the TTUdep for a $<$package, dependency$>$ relationship in Section 4.2. Formally speaking, TTUdep of a package $p_{i}$ with considering only dependency $p_{j}$ ($<p_{i},p_{j}>$ relationship) is:
(2) ${TTU}_{dep}(p_{i},p_{j})=\sum\operatorname{intervals\ with\ updated=False}$
MTTUdep of package $p_{i}$ with $n$ direct dependencies is:
(3) ${MTTU}_{dep}(p_{i})=\frac{\sum\limits_{j=1}^{n}{TTU}_{dep}(p_{i},p_{j})}{n}$
Metric 2 ($\operatorname{\mathbf{Mean-Time-To-Remediate}}_{dep}$ : $\mathbf{MTTR}_{dep}$).
MTTRdep of a package is the average aggregated time a package uses an outdated and vulnerable direct dependency version in its lifetime.
We have described how to calculate the TTRdep for a $<$package, dependency$>$ relationship in Section 4.2. Formally speaking, TTRdep of a package $p_{i}$ with considering only dependency $p_{j}$ ($<p_{i},p_{j}>$ relationship) is:
(4) ${TTR}_{dep}(p_{i},p_{j})=\sum\operatorname{intervals\ with\ remediated=False}$
MTTRdep of package $p_{i}$ with $n$ direct historically vulnerable dependencies is:
(5) ${MTTR}_{dep}(p_{i})=\frac{\sum\limits_{j=1}^{n}{TTR}_{dep}(p_{i},p_{j})}{n}$
5. Methodology
In this section, we first present vulnerability, package metadata (versions and dependency relations), and package characteristics collection process. We then present a statistical testing process to verify if MTTUdep can be a proxy for MTTRdep (RQ1). We then present correlation tests and regression analysis to understand how package characteristics impact MTTUdep and MTTRdep (RQ1).
5.1. Data Collection
5.1.1. Vulnerability Information
We collect the CVE data for our chosen three ecosystems from osv.dev (osv-dev, ) for npm, PyPI, and Cargo packages on 2024-09-12. We rely on OSV since OSV aggregates CVE data from multiple sources (e.g., GitHub Security Advisories, PyPA, GoVulDB) (osv-data-sources, ) in one place. After downloading JSON formatted CVE data from OSV, we convert it into an SQL table that includes ecosystem, package name, CVE ID, version where the vulnerability was introduced, and version where the vulnerability is fixed.
5.1.2. Package Metadata
We collect the package-version data and dependency information for npm, PyPI, and Cargo packages from deps.dev on 2024-08-20, similar to other previous studies (lin_untrustide_2024, ; zahan_software_2023, ). In our dataset, we have initially $2,603,314$ npm, $274,720$ PyPI, and $122,069$ Cargo packages. We collected data from deps.dev since it provides all the package versions and dependency information for our three chosen ecosystems. After data collection, we split each $<$package, dependency$>$ relation into multiple intervals. Since the dependency resolution can be complex and differs across ecosystems, we use deps.dev to perform the dependency resolution allowed by the dependency constraints with our added requirement of only using the available versions of the dependency before the interval start time.
5.1.3. Package Characteristics
We examined several characteristics (from Saini et al. (saini_investigating_2020, )) for packages: the number of contributors; the number of dependencies and dependents; the number of version releases; the SourceRank score, and the number of forks and stars. On 2025-01-11, we downloaded these characteristics of each package from libraries.io to understand if these factors correlate with package’s MTTUdep and MTTRdep values. Libraries.io is used by other research as a data source (gu_self-admitted_2023, ; cao_towards_2023, ; he_migrationadvisor_2021, ). When counting the number of dependencies, we only use the number of dependencies of packages in their latest version. To ensure construct validity in downloading these factors, we manually inspected a sample of packages to verify the factors, and we found the downloaded factors to be accurate. Additionally, we computed the number of major versions and package ages for our packages and added them to our analysis.
5.2. Package Inclusion Exclusion Criteria
We begin with an initial dataset of $3,000,103$ ($2,603,314$ npm, $274,720$ PyPI, and $122,069$ Cargo) packages collected from deps.dev. Our first step is to apply two inclusion criteria: (1) the package must be at least two years old (operationalized by the difference between the first and last version release); and (2) the package must have at least one residual activity (e.g., one version release) in the last two years. Miller et al. (miller_understanding_2025, ) used two years of residual activity followed by two years of no maintenance as criteria to find out the abandoned packages. Our criteria are inspired by Miller et al. since we want to include the packages that are maintained. Moreover, “two years” is a commonly used standard adopted by other research to measure whether a package is maintained or not (li_comparison_2023, ). However, our package selection criteria might miss packages that are less than two years old or have had no activity in the last two years (e.g., feature complete packages (coelho2017modern, )) even if they are not abandoned. We then used our exclusion criteria: a package without any dependency should be excluded. Since our metrics characterize packages’ dependency update practice, packages without any dependencies do not fit into our study. The number of packages after this exclusion criteria is $163,207$ ($117,129$ npm, $42,777$ PyPI, and $3,301$ Cargo packages). Out of these $163,207$ packages, $22,513$ ($17,263$ npm, $5,158$ PyPI, and $92$ Cargo) packages have at least one vulnerable dependency in their lifetime, supporting our initial motivation of lack of vulnerability data.
5.3. Metrics Implementation
After resolving the dependencies and applying our inclusion-exclusion criteria, we store the data in a PostgreSQL database. Given the large size of the dataset, working directly with the raw data would be impractical. We also create indexes on the most frequently accessed keys to speed up data retrieval for our metrics calculations.
In the database, we compute if the resolved dependency version matches the highest available version of the dependency during each interval and mark each interval accordingly, as shown in the “updated” column in Table. 2. Similarly, we then compute if the resolved dependency version for each interval is vulnerable to any security advisory even though a fixed version is available and mark the interval accordingly in the “remediated” column in Table. 2.
5.4. Statistical and Correlation Tests
To answer RQ1, we claim that if MTTUdep and MTTRdep are statistically equivalent, then MTTUdep can serve as a proxy for MTTRdep. Since we did not expect our samples to follow a normal distribution, we initially selected the non-parametric Mann-Whitney test. However, we also applied the t-test to ensure a more conservative analysis, as our data is not ordinal. Both tests were conducted using the Python library scipy, and the Wilcoxon test was also implemented using the same package. To assess normality, we used the Kolmogorov-Smirnov test.
Our initial hypotheses focused on the equivalence of data distributions. For these, following the approach of Papotti et al. (papotti_acceptance_2024, ), we employed the TOST (Two One-Sided Test) approach, originally proposed by Schuirmann et al. (schuirmann_hypothesis-testing_1981, ) and commonly used in pharmacological and food sciences to evaluate equivalence within a specified range (guidance_statistical_2001, ; meyners_equivalence_2012, ). Other research also used the TOST test in their analysis (shelton_rosita_2021, ; labunets2018no, ; labunets2017graphical, ; palheiros2024does, ). According to TOST, two distributions, $x$ and $y$, are considered equivalent if $x\cdot\delta<y<x\cdot 1/\delta$, where $\delta=0.8$. The TOST method can use either Mann-Whitney test or t-test as its underlying direction test. We reported the results of the t-test because it was more conservative than the Mann-Whitney test.
Additionally, to measure the association between MTTUdep and MTTRdep, we calculated both the Pearson correlation coefficient and Spearman’s rank correlation coefficient (similar to Cox et al. (cox_measuring_2015, )). High ($>0.7$) or moderate coefficient values ($>0.5$) would suggest a strong to moderate positive relationship (ratner2009correlation, ). Such findings would also support the notion that MTTUdep could effectively serve as a proxy for MTTRdep.
5.5. Package Characteristics Analysis
For analyzing eight package-level characteristics in RQ1, we use correlation tests to estimate if there is any correlation between these characteristics and packages’ corresponding MTTUdep and MTTRdep values. Specifically, we test a hypothesis about the association between the following independent variables and their continuous dependent variables (MTTUdep and MTTRdep):
Contributors Count ($H_{1}$). We hypothesize packages with fewer contributors are less likely to have updated dependencies since they have less capacity for maintenance and dependency management.
Dependents Count ($H_{2}$). We hypothesize packages with fewer dependents are less likely to have updated dependencies. Fewer dependents may lead to fewer feature updates, fewer bug fixes, and fewer version releases, which in turn may result in less updated dependencies.
Dependency Count ($H_{3}$). We hypothesize packages with fewer dependencies are more likely to have updated dependencies. We expect fewer dependencies to be more manageable, and thus, these projects may have more updated dependencies.
Version Count ($H_{4}$). We hypothesize packages with fewer number of version releases are less likely to have updated dependencies. Since these packages have less available versions, they may pay less attention to their dependency management.
Major Version Count ($H_{5}$). We hypothesize packages with fewer major version releases are more likely to have updated dependencies. Since these packages have fewer major versions to maintain, they may pay more attention to their dependency management.
Package Age ($H_{6}$). We hypothesize packages with lower ages are more likely to have updated dependencies. Lower package age may mean the development team is more proactive in dependency management since the package is not old or mature enough. Lower package age may also mean that the dependencies do not have scope for publishing many newer versions, which in turn may mean more updated dependencies in these packages.
SourceRank ($H_{7}$). The Package SourceRank score indicates the package quality, popularity, and community metrics calculated in libraries.io dataset (saini_investigating_2020, ; sun_using_2023, ). This metric depends on several factors, such as the presence of a README file, license, following SEMVER, recent updates, and the number of contributors. We hypothesize packages with lower SourceRank scores are less likely to have updated dependencies. Since these packages are of low quality, they may not have a lot of dependents or contributors, which in turn may result in less updated dependencies.
Forks And Stars Count ($H_{8}$). We hypothesize packages with fewer forks and stars are less likely to have updated dependencies. Fewer forks and stars may mean fewer people are using and looking into these packages, which in turn may mean less activity, less version releases, and less updated dependencies in these packages.
5.6. Regression Analysis
Correlation results indicate whether two independent variables are associated with each other or not. However, correlation does not establish causation. To identify such a cause-and-effect relationship using mathematical modeling, we opt for applying a technique of multilinear regression analysis. The multilinear regression model takes independent variables (e.g., package characteristics) and one dependent variable (MTTUdep or MTTRdep) and gives results on the relationship between each independent variable and the dependent variable. This analysis provides valuable insights on predicting MTTUdep or MTTRdep using the package’s characteristics. This analysis also results in $p$-values and coefficients, indicating which package characteristics might have a significant impact on MTTUdep or MTTRdep.
This process will produce a $p$-value for each independent variable to indicate whether the relationship between this variable and the outcome is statistically significant. Finally, we consider the variables with $p$-value $<0.05$ as significant variables, and the package characteristics associated with these variables are identified as key package characteristics. We use statsmodels package of Python to conduct the correlation tests and to build the regression model.
6. Results
6.1. RQ1: Can MTTUdep serve as a proxy for MTTRdep?
We statistically assess the relationship between MTTUdep and MTTRdep to determine if MTTUdep can effectively serve as a proxy for MTTRdep in  RQ1.
We used TOST (Two One-Sided Test) to investigate the equivalence of MTTUdep and MTTRdep. The single one-sided tests are done with the Mann-Whitney test. The results of the TOSTs are reported in Table 3. We also visualized the TOST for MTTUdep and MTTRdep from all three ecosystems in Fig. 2.
Table 3. TOSTs Results
Ecosystem MTTUdep ×$\times 0.8$ $<$ MTTRdep $<$ MTTUdep ÷$\div 0.8$
[…]
The green area corresponds to $x\cdot\delta<y<x\cdot 1/\delta$ where $\delta=0.8$, $x$ is MTTUdep, and $y$ is MTTRdep.
Figure 2. Distribution behind TOST MTTUdep vs. MTTRdep.
The results of the TOSTs show statistical equivalence between MTTUdep and MTTRdep ($p<0.05$) within the given range. Further, we also ran correlation analyses between MTTUdep and MTTRdep, in which both Spearman and Pearson correlation coefficients are $0.5$. The correlation analysis supports the results of our TOST tests as we can conclude that MTTUdep and MTTRdep are moderately and positively correlated since all correlation coefficients are $>=0.5$.
This result confirms that as the MTTUdep increases for a package, it is highly probable that the MTTRdep also increases (and vice versa), supporting the use of MTTUdep as a proxy for MTTRdep when we do not have enough vulnerability data to compute MTTRdep.
Takeaway 1: MTTUdep can be used as a proxy for MTTRdep since MTTUdep and MTTRdep are statistically equivalent (and moderately and positively correlated) in all three observed ecosystems.
6.2. RQ1: How do other package characteristics influence MTTUdep and MTTRdep?
Figure 3. Correlation matrix on packages’ characteristics and MTTUdep and MTTRdep values.
6.2.1. Correlation Analysis.
We tested eight package-level characteristics that might influence MTTUdep and MTTRdep. We present the correlation matrix with the packages’ characteristics and MTTUdep and MTTRdep data in Figure 3. The correlation heatmap provides insight into the relationships between various factors and their potential impact on MTTUdep and MTTRdep. The results indicate weak correlations between MTTUdep and the other factors.
Contributors Count ($H_{1}$): We hypothesized that packages with fewer contributors would be less likely to maintain updated dependencies. From the heatmap, we can see that the contributors count show minimal correlations with MTTUdep (-0.04) and MTTRdep (-0.07), suggesting a limited influence on dependency management.
Dependents Count ($H_{2}$): We hypothesized that packages with fewer dependents are less likely to have updated dependencies. The results indicate a negligible correlation between dependents count and MTTUdep or MTTRdep. This suggests that the number of dependents does not play a substantial role in determining the timeliness of dependency updates.
Dependency Count ($H_{3}$): We hypothesized packages with fewer dependencies to have more updated dependencies. The results show almost no correlation between dependency count and MTTUdep or MTTRdep. This implies that the number of dependencies in a package is not associated with dependency management efficiency.
Version Count ($H_{4}$): We hypothesized that packages with fewer version releases are less likely to have updated dependencies. However, the correlation matrix revealed no meaningful relationship between the number of version releases and either MTTUdep or MTTRdep. This suggests that the frequency of version releases does not significantly impact how quickly dependencies are updated.
Major Version Count ($H_{5}$): We hypothesized that packages with fewer major version releases are more likely to have updated dependencies. However, the correlation matrix revealed no meaningful relationship between the number of major version releases and either MTTUdep or MTTRdep. This suggests that the number of available major versions to maintain does not significantly impact how quickly dependencies are updated.
Package Age ($H_{6}$): We hypothesized that packages with lower ages are more likely to have updated dependencies. However, the correlation matrix invalidated this hypothesis with a weak correlation score ($0.32$). This suggests that as packages become mature with age, they tend to keep dependencies less updated. However, this conclusion should be treated with caution since our metrics might be age-sensitive (more discussion on this in Section 7).
SourceRank ($H_{7})$. We hypothesized that low quality packages (e.g., packages with lower SourceRank scores) are less likely to have updated dependencies. The correlation matrix supports this hypothesis for MTTRdep with a weak correlation score ($-0.22$), which means low-quality packages have higher MTTRdep values. However, low quality packages do not necessarily have higher MTTUdep values ($-0.06$ coefficient).
Forks And Stars Count ($H_{8}$): We hypothesized that packages with fewer forks and stars are less likely to have updated dependencies. The results reveal weak to moderate positive correlations between stars (0.65) and forks (0.72) with maintainers count, indicating that more popular packages may have more active development. However, their relationships with MTTUdep and MTTRdep remain insignificant, suggesting that forks and stars do not directly translate into faster dependency updates.
Overall, our results indicate weak or negligible correlations between the tested factors and dependency management efficiency as measured by MTTUdep and MTTRdep. These findings suggest that the analyzed factors may not strongly influence how updated dependencies are, and further exploration of other variables or non-linear relationships is recommended.
Takeaway 2: Package characteristics (except package age and SourceRank score) have negligible correlations with dependency updatedness. Older packages are associated with higher MTTUdep, and popular and better quality packages (higher SourceRank score) are associated with lower MTTRdep.
6.2.2. Regression Analysis (MTTUdep As Dependent Variable).
Interpreting the multilinear regression model’s characteristics, we found 0.073 as R-squared value when MTTUdep was the dependent variable. R-squared value 0.073 indicates that 7.3% variation in the dependent variable can be explained by the model. Although the model is statistically significant (p-value ¡ 0.05 from the F-test), lower R-squared value indicates that other factors, beyond the nine dependent variables, also substantially impact MTTUdep. The F-statistic of this model is large (1399), and Prob(F-statistic) is 0.00. This indicates that at least one of the independent variables (or predictors) has a non-zero relationship with MTTUdep.
We then look into the coefficients and p-values associated with each of the independent variables. The coefficient indicates the expected change in the dependent variable for a one-unit change in one independent variable while holding other independent variables constant. The p-value associated with the t-statistic indicates whether the independent variable is statistically significant in explaining the variation in the dependent variable. A lower p-value (¡ 0.05) would indicate statistical significance.
We have found that some independent variables have positive coefficients, which indicates an increase in one of these variables would result in an increase in our dependent variable, MTTUdep. Independent variables with positive coefficients are contributors count (0.0204), stars (0.0035), dependencies count (2.3027), and package age (0.1257). The dependencies count has the largest positive coefficient, which means that packages with more dependencies need longer times to update their dependencies. In addition, older packages tend to have longer MTTUdep as indicated by the 0.1257 coefficient, justifying our hypothesis ($H_{3}$) on dependencies count.
We also found some independent variables have negative coefficients. Negative coefficients indicate that an increase in that independent variable would result in a decrease in the dependent variable, MTTUdep. Independent variables with negative coefficients are dependents count (-0.0013), forks (-0.0085), SourceRank (-24.3203), and major versions count (-0.9068). SourceRank, which indicates package quality and popularity, shows strongly negative coefficients. Higher SourceRank indicates an overall better project. The model predicts if SourceRank for a package increases by 1 (moving to a “better” rank), MTTUdep drops by 24 days. Moreover, packages that have released more major versions tend to have a lower MTTUdep. Finally, all dependent variables except dependent count were found to be statistically significant (p-value ¡ 0.05).
Takeaway 3: Packages with a lot of dependencies or older packages are more likely to have higher MTTUdep. Also, popular and better quality packages (with higher SourceRank scores) are more likely to have lower MTTUdep.
6.2.3. Regression Analysis (MTTRdep As Dependent Variable).
For MTTRdep as the dependent variable, the multilinear regression model results in 0.097 as R-squared value. R-squared = 0.097 indicates that the model explains 9.7% of the variation in MTTRdep. A lower R-squared value indicates that other unmodeled factors play an important role in determining MTTRdep, which is similar to our observations in MTTUdep. The F-statistic of this model is 246.8, and Prob(F-statistic) is 0.00. This indicates the model is statistically significant and at least one of the dependent variables (or predictors) has a non-zero relationship with MTTRdep.
We then look into the coefficients and p-values associated with each of the independent variables. We found that forks, stars, SourceRank, dependencies count, major version count and package age are statistically significant (p-value ¡ 0.05). On the other hand, the contributors count, and dependents count are not statistically significant (p-value ¿ 0.05) in modeling MTTRdep.
As a predictor, forks (coefficient 0.0070) and stars (coefficient 0.0014) indicate a positive but small effect. More forks and stars are associated with slightly higher MTTRdep. Dependencies count (coefficient 0.5028), and package age (coefficient 0.1239) show a moderate positive effect on MTTRdep. Additionally, the major version count (coefficient -2.4954) shows a moderate negative effect, whereas SourceRank (coefficient -34.8183) shows a strong negative effect on MTTRdep. This observation is similar to our previous observation with MTTUdep. According to the model’s prediction, an increase in SourceRank by 1 (i.e., a higher SourceRank indicates a better package) would result in a 34.8-day reduction in MTTRdep.
Takeaway 4: Similar to MTTUdep, older packages are more likely to have higher MTTRdep. Also, popular and better quality packages (with higher SourceRank scores) are more likely to have lower MTTRdep. Packages with higher major versions are more likely to have lower MTTRdep.
6.3. RQ1: How long do packages have outdated dependencies, and how long do packages have vulnerable dependencies?
In RQ1, we empirically analyze the MTTUdep and MTTRdep metrics for the three ecosystems. We choose violin plot instead of box-plot since violin plot shows everything box-plot shows, e.g., medians, ranges, variability, and violin plot’s shape shows the density of the data similar to a density estimation plot (hintzeViolinPlotsBox1998, ).
Figure 4. ${MTTU}_{dep}$
Figure 5. ${MTTR}_{dep}$
Fig. 4 shows a violin plot of MTTUdep in npm, PyPI, and Cargo. All the plots are right-skewed, indicating that most packages have a low MTTUdep. For instance, $50\%$ npm package has MTTUdep of less than $174$ days. In terms of distribution width, Cargo has a compact distribution, and npm is the most widely spread (Cargo ¡ PyPI ¡ npm). Also, every plot has a long tail, indicating that every ecosystem has some packages that do not update their dependencies for a long time (max MTTUdep $=4296$ days).
Fig. 5 shows a violin plot visualizing the distribution across MTTRdep in days for Cargo, npm, and PyPI packages. The first, second (mean), and third quantiles are shown as , , and , respectively. The overall pattern is similar to MTTUdep. The relative distribution width and the right-skewed nature of MTTRdep is similar to MTTUdep. Cargo and PyPI have concentrated distributions, with most MTTRdep (e.g., interquartile range) being less than $219$ and $281$ days. Conversely, npm shows a more spread distribution than Cargo and PyPI, indicating variability in MTTRdep in npm.
We could compute MTTUdep for $163,207$ ($117,129$ npm, $42,777$ PyPI, and $3,301$ Cargo) packages, and MTTRdep for $22,513$ ($17,263$ npm, $5,158$ PyPI, and $92$ Cargo) packages. This corroborates our initial motivation for conducting this study, the lack of vulnerability data.
Takeaway 5: Even without potentially abandoned packages, a subset of packages show extreme delays in keeping their dependencies and vulnerable dependencies updated.
Comparison to Prior Work. Similar to prior research (kula_developers_2018, ), we also found that the majority of packages in an ecosystem do not keep dependencies updated. We also observed a long-tail distribution for both MTTUdep and MTTRdep in all three ecosystems. A significant number of packages lag behind in keeping dependencies updated, which is similar to the observations of Cox et al. (cox_measuring_2015, ). In short, regularly updating dependencies is not a widespread practice. Our observation is similar to vulnerable dependency updates as well. Many packages took months or longer to remediate known vulnerable dependencies, similar to observations from Kula et al. (kula_developers_2018, ). This suggests that challenges in keeping dependencies updated are not specific to any ecosystem but rather general in nature.
7. Discussions And Implications
When MTTUdep and MTTRdep should be used. Our study provides new insights by establishing the relationship between MTTUdep and MTTRdep. We found that although a package’s MTTUdep is moderately correlated with its MTTRdep, MTTUdep and MTTRdep are statistically equivalence. In practical terms, packages that are slow to update dependencies also tend to be slow in updating vulnerable dependencies. This finding implies that improving general dependency update practice will likely improve vulnerable dependency update practice, which seems logical, but we provide statistical evidence supporting our hypothesis. Having said that, developers should use MTTRdep when available and only MTTUdep as a proxy for MTTRdep when MTTRdep is not available.
Package Age’s Influence In MTTUdep and MTTRdep. Package age is weakly correlated with MTTUdep and has a negligible correlation with MTTRdep in our analysis. In our regression analysis, we have used package age as an independent variable with MTTUdep and MTTRdep as dependent variables. However, since we design our metrics based on a temporal perspective, our metrics could have an inherent age sensitivity. So, conclusions like ‘older packages tend to have slightly higher MTTUdep’ should be treated cautiously. Since older packages might have more data points (negligible but positive correlation with version count, major version count, and dependencies count), package age could influence our analysis. Future research should explore how to model the package age in MTTUdep and MTTRdep (e.g., one way could be to normalize the metrics by package age or by only taking the last $n$ years into account).
Influence of Unmodeled Package Characteristics. Prior work shows that developers hesitate to update dependencies due to a fear of breaking changes, a lack of awareness or knowledge about available updates, and sometimes a lack of motivation to invest time in updates (derrKeepMeUpdated2017, ). These human factors are the likely reasons behind the long tail of MTTUdep and MTTRdep data. Our findings also support this. Packages with many contributors or higher dependents count do not have a better MTTUdep or MTTRdep, in contrast to our hypothesis. A package used by thousands of other packages is not guaranteed to have dependencies updated. Both multilinear regression models in our study present an R-squared value of less than 10%. A lower R-squared value indicates that other unmodeled factors substantially affect MTTUdep and MTTRdep. The development team dynamic, the use of pinning and floating, the cost and efforts needed in testing, and the size of the codebase could be such possible unmodeled factors.
Practical implications for developers. We confirmed that improving general dependency update frequency likely improves security as well. This should be a strong incentive for the developers to treat dependency updates as an important part of routine maintenance, not an afterthought. Even scheduling periodic dependency update sprints or using automated update notifications to stay on top of the new releases can be an effective strategy. Allowing the latest version of the dependency using the dependency specification (e.g., * or $latest$) is the best way to keep MTTUdep and MTTRdep low (even zero). However, that might not be possible due to the issues with breaking changes (ochoa_breaking_2022, ; raemaekers_semantic_2014, ; venturini_i_2023, ). Allowing auto updates of minor and patch releases by the dependency could help reduce the MTTUdep and MTTRdep. This ensures developers get incremental improvements and fixes without manual effort. Similarly, pinning should be avoided since pinning does not allow any auto-updates, which will start increasing the MTTUdep as soon as the dependency releases a new version (even a patch release). Whenever a vulnerability is found in the dependency and a fixed version is available, developers should prioritize remediating that vulnerability. Finally, a mixed strategy with automatic minor and patch updates, alongside manual major version updates, could be the most effective strategy to keep the MTTUdep and MTTRdep minimal. Our recommendation relies on developers using SEMVER correctly (pinckney_large_2023, ; li_large-scale_2023, ; dietrich_dependency_2019, ) so that the benefits of floating versions can be leveraged.
Gaming Metrics. While it might seem that using a loose floating constraint and bumping dependency constraint periodically could “game” our metrics, the underlying outcome is that a package spends minimal time with outdated dependency versions and ensures that security fixes are adopted automatically. By doing so, this strategy reduces the risk of exploitation from using an outdated dependency version with known security vulnerability, such as log4Shell (everson_log4shell_2022, ; rahman_less_2024, ). In short, trying to game our metrics by adopting floating version constraints is essentially reducing this attack vector for a package. However, this approach also comes with a tradeoff. This strategy might make the packages susceptible to malicious package updates (xz-incident (williams_research_2025, )). Since floating version constraints allow automatic updates, it can also facilitate propagating malicious package updates.
Influence of Dependencies Count in MTTUdep and MTTRdep. In our design, we compute MTTUdep and MTTRdep by dividing the summation of TTUdep and TTRdep by the number of dependencies and number of historically vulnerable dependencies, respectively, in Eq. 3 and  5. A package with a high number of dependencies has more room for having a better value in MTTUdep while keeping the majority of dependencies updated. However, MTTRdep is not influenced by dependencies count since we only consider the amount of time spent with vulnerable dependency versions and the number of historically vulnerable dependencies.
8. Threats to Validity
In this section, we discuss potential limitations and how they might impact the interpretation of our results.
External Validity. The main external validity threat is the generalizability of our results to characterize other ecosystems. We carefully chose the three ecosystems to have a diverse set of ecosystems, where npm is the largest, PyPI is the oldest, and Cargo is the newest among the major software ecosystems. While each ecosystem possesses unique features that might not directly correlate with those we studied, we believe the insights gained should also be broadly applicable to other ecosystems.
Internal Validity. We use the security advisory dataset from OSV.dev, which may not be comprehensive. If an advisory is published but not included in the OSV dataset, that may impact our results. Additionally, we do not consider whether the vulnerable dependency version is exploitable or reachable (cisa-vex, ) from the package. We treat all vulnerabilities equally regardless of the CVSS score or the severity of the vulnerability. After downloading the data from deps.dev, we manually checked 20 packages’ versions and relations with the public package registries and found that the data is accurate. Moreover, each package manager has its own way of handling dependency resolutions, and for the dependency resolutions, we rely on the Open Source Insights (depsdev, ) resolved version data. Our analysis omits package versions not adhering to SEMVER rules, a conservative choice to enable a more rigorous analysis. In addition, Open Source Insights dependency resolution fails in some cases (e.g. missing timestamp), and we mark those as warnings in our dataset. We do not calculate update metrics for those packages, and we argue that this might have a very small impact on our results. In addition, we only consider runtime dependencies in our MTTUdep and MTTRdep analysis and omit dev and optional dependencies. Additionally, some dependencies might be more important than others depending on the context; however, we treat each dependency equally since modeling dependencies’ importance is out of the scope of our study.
9. Conclusion and Future Works
In this study, we introduced two dependency update metrics, MTTUdep and MTTRdep, to quantify the updatedness of dependencies in open-source software packages. Our large-scale empirical analysis across the npm, PyPI, and Cargo ecosystems demonstrated that MTTUdep can serve as a proxy for MTTRdep when vulnerability information is unavailable. Furthermore, our statistical analysis highlighted the relationships between package characteristics and dependency update behavior, providing actionable insights for developers, maintainers, and software supply chain researchers. Future research can explore additional factors influencing dependency update practices, such as the severity of vulnerabilities, organizational policies, or developer incentives. Expanding this analysis to other ecosystems, combining with transitive dependencies, and incorporating qualitative insights from developers could further refine our understanding of dependency updatedness.